OpenStudy (archtangent):

can anyone tell me how "Dynamics of Rigid Bodies" is connected to electronics engr

OpenStudy (archtangent):

halp

OpenStudy (sidsiddhartha):

dynamics of rigid bodies are mainly connected with control systems which is a part of electronics engineering. you can form a mathematical model of the dynamic system by writing diffrential equations and then convert it to a equivalent electronics/electrical network . take this as an example-- \[M \frac{ d^2x }{ dt^2 }+B.\frac{ dx }{ dt }+Kx=F(t)..........(1)\] this is a equation of a dynamic system now u can compare it with equation of a simple RLC series circuit--\[v(t)=R.i+L.\frac{ di }{ dt }+\frac{ 1 }{ c }\int\limits_{0}^{t}i.dt\\which ~is~equivalent ~\to\\v(t)=L.\frac{ d^2q }{ dt^2 }+R.\frac{ dq }{ dt }+\frac{ q }{ C }................(2)\] now compare between these two equations and u'll get \[F(t) \rightarrow V(t)\\x \rightarrow q\\k \rightarrow \frac{ 1 }{ C}\\B \rightarrow R\\M \rightarrow L\\\] so you can see that any dynamics system can be converted into a electrical or a electronic system ,i think that is the connection you are looking for @ArchTangent

OpenStudy (kenljw):

What was demonstrated is what used to be called a electric analog computer, but I don't think that is what your asking. There are also mechanical analog configuration the must be considered. The use of rigid bodies and electrical is used in every motor and generator, and in my day people got Electrical Mechanical degrees to come to understand both system and see how the correlated.